Based on a poll 60% of adults believe in reincarnation. Assume that 4 adults are randomly selected, and find the indicated probability.


1
Expert's answer
2021-09-27T18:11:43-0400

a) "P(X=3)=C_4^30.6^40.4^1=0.346."

b) "P(X=4)=C_4^40.6^40.4^0=0.130."

c) "P(X\ge3)=P(X=3)+P(X=4)=0.475."